OpenStudy (anonymous):

Find the slope of line passing through the points (4,8) and (6, -2). A. 1/5 B. 5 C. -1/5 D. -5

jimthompson5910 (jim_thompson5910):

m = (y2-y1)/(x2-x1) m = ( -2-8)/(6-4) m = -10/2 m = -5 So the slope of the line that passes through the two points is -5

OpenStudy (anonymous):

You use the slope formula: \[m=\frac{y _{1}-y _{2}}{x _{1}-x_{2}}\] and plug in the points given to you: \[m=\frac{8-(-2)}{4-6}\] Simplify and get that \[m=\frac{10}{-2} or -5\] So your answer is D :)